OpenStudy (anonymous):

For each expression write an equivalent expression without parentheses: 4(3x+5) 4(3x-5) -4(3x+5) -4(3x-5)

OpenStudy (mysangels):

You have to distribute to make an equivalent expression by using the distributive property

OpenStudy (anonymous):

What happens if you just give me the answer? Do you get kicked off of open study or something? Lol just wondering.

OpenStudy (mysangels):

-4(3x-5) -4 x 3x= -12x -4 x -5= 20 -12x + 20 is your finally answer now use the distributive property with the rest of the problems

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Are those all four answer to the four questions?

OpenStudy (solomonzelman):

I am going to write a rules you shall apply. 1) \(\large\color{slate}{ \color{red}{a}(\color{blue}{b}+\color{green}{c})= \color{red}{a} \cdot \color{blue}{b} ~+~\color{red}{a} \cdot \color{green}{c} }\) 2) \(\large\color{slate}{ \color{red}{a}(\color{blue}{b}-\color{green}{c})= \color{red}{a} \cdot \color{blue}{b} ~-~\color{red}{a} \cdot \color{green}{c} }\)

OpenStudy (mysangels):

In this case, yes.

OpenStudy (solomonzelman):

in any case. (if you are referring to my rules)

OpenStudy (mysangels):

these rules are very helpful if you have to distribute anything